# Peer-to-Peer Basis Risk Management for Renewable Production Parametric Insurance

### v1

@@ Abstract
This work presents a framework for peer-to-peer (P2P) basis risk management applied to solar electricity generation. The approach leverages physically based simulation models to estimate the day-ahead production forecasts and the actual realized production at the solar farm level. We quantify the financial loss from mismatches between forecasted and actual production using the outputs of these simulations. The framework then implements a parametric insurance mechanism to mitigate these financial losses and combines it with a P2P market structure to enhance participant risk sharing. By integrating day-ahead forecasts and actual production data with physical modeling, this method provides a comprehensive solution to manage production variability, offering practical insights for improving financial resilience in renewable energy systems. The results highlight the potential of combining parametric insurance with P2P mechanisms to foster reliability and collaboration in renewable energy markets.
@@
[…]
[fallou]organization=Centre de Recherche en Économie et de Statistiques (CREST), CNRS, École Polytechnique, GENES, ENSAE Paris, Institut Polytechnique de Paris, addressline=5 Avenue Henry Le Chatelier, city=Palaiseau, postcode=91120, country=France
[…]
[alicia]organization=Laboratoire de Génie Industriel (LGI), CentraleSupélec, addressline=3 Rue Joliot Curie, city=Gif-Sur-Yvette, postcode=91190, country=France
[…]
[michel]organization=Institut de statistique, biostatistique et sciences actuarielles - ISBA, Université Catholique de Louvain, addressline=1 Place de l’université, city=Louvain, postcode=B-1348, country=Belgium
[…]
[christian]organization=Institut de Science Financière et d’Assurances (ISFA), Université Claude-Bernard Lyon 1, addressline=50 Av. Tony Garnier, city=Lyon, postcode=69007, country=France
@@ 1 Introduction
The increasing reliance of renewable energy producers on market revenues introduces a significant source of uncertainty, which can deter investment decisions due to unpredictable cash flows. Several European countries have implemented Capacity Remuneration Mechanisms (CRMs) to address this, ensuring revenues are not solely tied to production, see e.g. Keppler (2017).
More recently, Contracts for Difference (CfDs) have emerged as a key instrument for stabilizing revenues, see e.g. Onifade (2016). Under a CfD, renewable energy producers are guaranteed a fixed “strike price” for their electricity. When market prices fall below this strike price, producers receive compensatory payments; conversely, when market prices exceed the strike price, producers return the surplus. This mechanism protects against price volatility, making renewable energy investments more predictable and attractive. However, while CfDs mitigate price risks, they fail to address production variability, leaving producers vulnerable to weather-related uncertainties and forecast errors. In addition to these market-based instruments, many European governments have introduced subsidies, such as feed-in tariffs and feed-in premiums, to stabilize renewable revenues. These policies provide guaranteed prices or additional premiums for renewable generation, ensuring predictable cash flows. However, they still do not eliminate the physical production risks.
@@
Parametric insurance has emerged as a promising mechanism for mitigating financial risks associated with weather-related variability in renewable energy production. Unlike traditional insurance, which compensates for actual losses, parametric insurance triggers payouts based on predefined parameters, such as weather indices. Han et al. (2019) explore weather index insurance for wind energy, showing how such contracts protect producers from economic losses due to unfavorable wind conditions by linking payments to weather indices rather than direct financial losses. Similarly, Liao et al. (2021) develop index-based insurance models using time-series forecasting historical wind speed data, building index insurance models to stabilize revenue and enhance investment security. Extending this approach to solar farms, Liao et al. (2022) design an index-based insurance product for photovoltaic plants in Taiwan, using long-term solar irradiation data and ARIMA models to calculate premiums and hedge against production shortfalls. The appeal of parametric insurance lies in its several advantages: it enables rapid payouts, relies on measurable third-party data (e.g., solar radiation or wind speed), and minimizes disputes by eliminating subjective claims assessments and human error. Moreover, by linking insurance triggers directly to measurable weather variables, insurers can more effectively hedge their exposure to weather-related losses in financial markets using weather derivatives. This enhances the stability of their portfolios and facilitates access to reinsurance and alternative risk transfer mechanisms, ultimately improving capital efficiency and enabling more sustainable coverage for renewable energy producers.
However, a key limitation of parametric insurance is basis risk—the mismatch between payouts and actual losses. Basis risk arises when the chosen parameters fail to capture the actual damages experienced, reducing the effectiveness of this mechanism, particularly in localized or atypical weather conditions. The issue of basis risk has been extensively discussed in the literature, with several solutions proposed. Figueiredo et al. (2018) advocates using probabilistic triggers, which incorporate uncertainty into the payout mechanism to better align compensation with actual losses. By accounting for variability in the relationship between weather parameters and financial damages, these triggers aim to reduce discrepancies contributing to basis risk. Despite these advancements, current approaches often overlook the potential for collective mechanisms to redistribute residual risks and improve participant resilience against weather variations.
@@
To address these limitations, we propose a novel hybrid model that combines parametric insurance with a P2P mechanism. In this context of P2P risk-sharing mechanisms, the study by Wang et al. (2025) provides valuable insights into the application of outage risk sharing in electric power systems. Their work highlights the benefits of decentralized risk-sharing frameworks in mitigating financial losses associated with power outages. This study strengthens our proposal by demonstrating how P2P mechanisms can be effectively implemented in the energy sector to enhance financial resilience. While Wang et al. (2025) focus on outage risks, our approach extends the concept to revenue volatility in renewable energy production. The P2P layer complements parametric insurance by redistributing residual risks among participants, providing an additional buffer accounting for basis risk. This approach mitigates individual exposure to financial losses caused by forecast errors and weather variability, enhancing the robustness of the insurance framework. By allowing producers to pool their risks, our model helps smooth out compensation mismatches and provides a more effective hedge against production uncertainties.
@@
P2P insurance, as implemented in our model, directly addresses the basis risk by redistributing financial impacts not covered by parametric insurance. Its main advantages include cost efficiency, as it eliminates overhead costs associated with traditional insurers; flexibility, as it can be tailored to meet the specific needs of renewable energy producers; and its cooperative structure, which strengthens risk-sharing among producers and aligns with the growing role of local energy initiatives in decentralized electricity markets. Our approach offers a practical and scalable solution for managing financial risks in renewable energy markets.
@@
The theoretical foundations of risk pooling further strengthen our approach. A key aspect is to determine the risk-sharing rule that has to be accepted by all producers. In this paper, we work with a risk-sharing rule based on an approximation of the conditional mean risk-sharing rule proposed in Denuit and Dhaene (2012). According to the conditional mean risk-sharing rule, each producer contributes the conditional expectation of the basis risk brought to the pool, given the total basis risk experienced by the entire pool (see Denuit and Robert (2021a)). Some attractive properties of the conditional mean risk allocation have been studied in Denuit (2019); Denuit and Robert (2020); Denuit et al. (2022); Jiao et al. (2022). As the number of participants in the pool grows, the conditional mean risk-sharing rule can be approximated by a linear risk-sharing rule. This principle is particularly relevant to our model, ensuring a simplified basis risk-sharing for large pools.
@@
Using solar energy in southern Germany as a case study, we simulate day-ahead forecasts and actual production at the solar farm level using a physics-based model and high-resolution climate data. From these simulations, we compute financial losses as the difference between anticipated and actual revenues. Using a reference location, we construct a weather-based index for parametric insurance and implement a P2P mechanism to redistribute residual basis risk among producers. Our results show that the P2P mechanism effectively mitigates financial losses associated with forecast errors and enhances revenue stability, reducing the volatility of losses by 55%. This improvement strengthens the financial viability of renewable energy investments, making them more resilient to weather-driven uncertainties.
@@
2 Model
In this section, we introduce the combined parametric and P2P insurance model.
@@ 2.1.1 Individual loss and basis risk
In renewable energy production, the random loss $X_{i}$ for producer $i$ is the financial loss associated with forecast uncertainties at a given date. The parametric insurance is triggered when an observed index $Z$ based on weather-related variables for a reference location (see Figure 1) reaches a predefined threshold $z_{0}$. Like the Consumer Price Index (CPI), which tracks changes in the cost of a standard basket of goods and services over time, the weather index $Z$ quantifies weather risks for renewable energy production through a composite of weather variables. An increase in the index reflects a rise in extreme weather events, leading to very low production, highlighting elevated risk levels for renewable energy systems.
@@ 2.1.2 Risk pooling and sharing
where $\sigma_{i}^{2}(z)=\text{Var}[X_{i}\mid Z=z],$ is the conditional variance of $X_{i}$ for producer $i$ given that the weather index $Z$ is equal to $z$ that exceeds the threshold $z_{0}$. This formula is not only a proportional allocation based on conditional variances but also serves as an approximation of the conditional expectation $\mathbb{E}[\varepsilon_{i}\mid S_{\varepsilon},Z=z]$ when $n$ is large, as shown in Denuit and Robert (2021b). It ensures fairness and aligns each producer’s contribution with the variability of their losses. The derivation of this formula as an approximation of $\mathbb{E}[\varepsilon_{i}\mid S_{\varepsilon},Z=z]$ assumes that the basis risks $\varepsilon_{i}$ are independent, conditional on $Z$. This conditional independence is crucial, as it requires the weather index $Z$ to act as a sufficiently informative linear score, capturing all potential conditional dependencies between individual losses. $Z$ should comprehensively reflect the shared weather-related risk factors influencing producers’ financial outcomes. By carefully designing $Z$, this assumption becomes a valid and practical simplification, ensuring the effectiveness and fairness of the risk-sharing mechanism.
@@ 2.2 Parameterizing the insurer’s index benchmark for effective risk sharing
In this paper, the construction of the index $Z$ draws inspiration from the North American Actuaries Climate Index (ACI) in the sense that both use a linear combination of climate variables (Garrido et al. (2024)). While the methodology used to build our index differs from the ACI, this linear structure is standard across many actuarial climate indices, providing an intuitive way to summarize weather conditions relevant to risk assessment and insurance design.
Remark 1
@@
We now detail the approximation of the following quantities of interest: $\mathbb{E}\left[X_{i}\mid Z=z\right]$ and $\text{Var}\left[X_{i}\mid Z=z\right]$.
@@ 2.2.1 Approximation of the conditional expectation $\mathbb{E}\left[X_{i}\mid Z=z\right]$
The terms $L_{i}(\cdot,\cdot)$ and $F_{i}(\cdot,\cdot)$ capture first- and second-order deviations, respectively, and adjust for individual differences in the weight vectors $\bm{a}_{i}$.
@@
The proof of the approximation of the conditional expectation of the loss in Equation (3) is given in Appendix B.
@@ 2.2.3 Final optimisation
3 Case study/Application
@@ 3.1.1 Power plants data
We compiled a dataset of large-scale solar power plants in Germany using the German governmental database ; see, e.g., Bundesnetzagentur (2024). This dataset was constructed by extracting and processing information on operational solar farms, focusing on installations exceeding 5 megawatt (MW). Such large-scale projects represent potential candidates for parametric insurance, as they primarily operate within electricity markets rather than for self-consumption.
@@ 3.1.2 Recovery of the energy production loss
To quantify these imbalances, we reconstruct both forecasted and realized conditions using meteorological variables from two datasets provided by the European Centre for Medium-Range Weather Forecasts (ECMWF). The historical weather variables are obtained from the ERA5 reanalysis dataset (Hersbach et al., 2020), which provides global coverage at a spatial resolution of 31 km with an hourly temporal resolution. This dataset, widely used in renewable energy research, spans 2012 to 2022 and includes near-surface air temperature (T2m) and surface solar radiation downwards (SSRD), which are essential for forecasting solar production.
@@
For each given location, we model both forecasted and realized solar electricity production using the PVlib library (Holmgren et al., 2018) in Python, which simulates photovoltaic system performance based on climate variables. The energy production loss for each solar farm is then defined as the difference between its forecasted and actual production. The PVlib model relies on several key inputs, including SSRD, T2m, panel orientation, and panel tilt angle. The tilt angle, which represents the inclination of the solar panels relative to the horizontal plane, is set at 25° in our simulations. This value corresponds to the mean tilt across our dataset, where individual panel angles range from 20° to 30°.
To recover energy output, we first compute Direct Normal Irradiance (DNI) and Diffuse Horizontal Irradiance (DHI) from SSRD using the Erbs model (Erbs et al., 1982). Since these values are not directly available, they must be derived using standard transposition models. The SSRD is then converted into Plane of Array (POA) irradiance, representing the amount of solar radiation incident on the tilted panel surface. The module temperature is estimated using a thermal model that accounts for both air temperature and POA irradiance. Finally, the PVWatts model (Dobos, 2014) is used to compute direct current (DC) power, which is then converted into alternating current (AC) power using an inverter model with default efficiency parameters.

### v2

@@ Abstract
The financial viability of renewable energy projects is challenged by the variability and unpredictability of production due to weather fluctuations. This paper proposes a novel risk management framework combining parametric insurance and peer-to-peer (P2P) risk sharing to address production uncertainty in solar electricity generation. We first design a weather-based parametric insurance scheme to protect against forecast errors, recalibrated at the site level to mitigate geographical basis risk. To handle residual mismatches between insurance payouts and actual losses, we introduce a complementary P2P mechanism that redistributes the remaining basis risk among participants. The method leverages physically based simulation models to reconstruct day-ahead forecasts and realized productions, integrating climate data and solar farm characteristics. A second-order theoretical approximation links heterogeneous local models to a shared weather index, making risk sharing operationally feasible. In an empirical application to 50 German solar farms, our approach reduces the volatility of production losses by 55%, demonstrating its potential to stabilize revenues and strengthen the resilience of renewable investments.
@@ 1 Introduction
The increasing reliance of renewable energy producers on market revenues introduces a significant source of uncertainty, which can deter investment decisions due to unpredictable cash flows. Power Purchase Agreements (PPA) were among the first solutions proposed to mitigate this uncertainty [1]. Since then, several European countries have implemented Capacity Remuneration Mechanisms (CRMs) to address this, ensuring revenues are not solely tied to production, see e.g. Keppler [2].
More recently, Contracts for Difference (CfDs) have emerged as a key instrument for stabilizing revenues, see e.g. Onifade [3]. Under a CfD, renewable energy producers are guaranteed a fixed “strike price” for their electricity. When market prices fall below this strike price, producers receive compensatory payments; conversely, when market prices exceed the strike price, producers return the surplus. This mechanism protects against price volatility, making renewable energy investments more predictable and attractive. However, while CfDs mitigate price risks, they fail to address production variability, leaving producers vulnerable to weather-related uncertainties and forecast errors. In addition to these market-based instruments, many European governments have introduced subsidies, such as feed-in tariffs and feed-in premiums, to stabilize renewable revenues. These policies provide guaranteed prices or additional premiums for renewable generation, ensuring predictable cash flows. However, they still do not eliminate the physical production risks.
@@
Parametric insurance has emerged as a promising mechanism for mitigating financial risks associated with weather-related variability in renewable energy production. Unlike traditional insurance, which compensates for actual losses, parametric insurance triggers payouts based on predefined parameters, such as weather indices. Han et al. [4] explore weather index insurance for wind energy, showing how such contracts protect producers from economic losses due to unfavorable wind conditions by linking payments to weather indices rather than direct financial losses. Similarly, Liao et al. [5] develop index-based insurance models using time-series forecasting historical wind speed data, building index insurance models to stabilize revenue and enhance investment security. Extending this approach to solar farms, Liao et al. [6] design an index-based insurance product for photovoltaic plants in Taiwan, using long-term solar irradiation data and ARIMA models to calculate premiums and hedge against production shortfalls. The appeal of parametric insurance lies in its several advantages: it enables rapid payouts, relies on measurable third-party data (e.g., solar radiation or wind speed), and minimizes disputes by eliminating subjective claims assessments and human error. Moreover, by linking insurance triggers directly to measurable weather variables, insurers can more effectively hedge their exposure to weather-related losses in financial markets using weather derivatives. This enhances the stability of their portfolios and facilitates access to reinsurance and alternative risk transfer mechanisms, ultimately improving capital efficiency and enabling more sustainable coverage for renewable energy producers.
However, a key limitation of parametric insurance is basis risk—the mismatch between index-based payouts and actual losses when the chosen parameters fail to capture local damages. In particular, a component of this basis risk, known as geographical basis risk, arises when the weather index is computed at a single reference station whose conditions diverge from those at individual farms, undermining the insurance’s effectiveness in atypical or micro-climatic settings [7]. The literature has proposed various solutions—e.g. probabilistic triggers that introduce uncertainty into the payout rule to better align compensation with realized losses [8]—yet these approaches still leave residual risks unaddressed. To mitigate geographic basis risk ex-ante, we therefore recalibrate the trigger for each site by fitting a simple farm-level model linking reference-index deviations to that farm’s historical production losses. This preliminary step smooths out spatial idiosyncrasies and significantly reduces basis risk.
@@
Despite site-specific recalibration, some residual mismatches inevitably persist. To address these remaining gaps, we layer a peer-to-peer (P2P) basis risk sharing mechanism on top of our recalibrated parametric contracts. In this context of P2P risk-sharing mechanisms, the study by Wang et al. [9] provides valuable insights into the application of outage risk sharing in electric power systems. Their work highlights the benefits of decentralized risk-sharing frameworks in mitigating financial losses associated with power outages. A complementary line of research by Moret et al. [10] has already shown that pure P2P risk‐sharing, when coupled with simple financial contracts in a community‐based electricity market, can significantly attenuate participants’ risk aversion and restore fairness in cost allocation. This study strengthens our proposal by demonstrating how P2P mechanisms can be effectively implemented in the energy sector to enhance financial resilience. While Wang et al. [9] focus on outage risks, our approach extends the concept to revenue volatility in renewable energy production. The P2P layer complements parametric insurance by redistributing residual risks among participants, providing an additional buffer accounting for basis risk. This approach mitigates individual exposure to financial losses caused by forecast errors and weather variability, enhancing the robustness of the insurance framework. By allowing producers to pool their risks, our model helps smooth out compensation mismatches and provides a more effective hedge against production uncertainties.
@@
P2P insurance, as implemented in our model, directly addresses the basis risk residual by redistributing financial impacts not covered by parametric insurance. Its main advantages include cost efficiency, as it eliminates overhead costs associated with traditional insurers; flexibility, as it can be tailored to meet the specific needs of renewable energy producers; and its cooperative structure, which strengthens risk-sharing among producers and aligns with the growing role of local energy initiatives in decentralized electricity markets. Our approach offers a practical and scalable solution for managing financial risks in renewable energy markets.
@@
The theoretical foundations of risk pooling further strengthen our approach. A key aspect is to determine the risk-sharing rule that has to be accepted by all producers, a question addressed frequently [11]. In this paper, we work with a risk-sharing rule based on an approximation of the conditional mean risk-sharing rule proposed in Denuit and Dhaene [12]. According to the conditional mean risk-sharing rule, each producer contributes the conditional expectation of the basis risk brought to the pool, given the total basis risk experienced by the entire pool (see Denuit and Robert [13]). Some attractive properties of the conditional mean risk allocation have been studied in Denuit [14], Denuit and Robert [15], Denuit et al. [16], Jiao et al. [17]. As the number of participants in the pool grows, the conditional mean risk-sharing rule can be approximated by a linear risk-sharing rule. This principle is particularly relevant to our model, ensuring a simplified basis risk-sharing for large pools.
@@
Main contributions
On the theoretical side, we derive a novel second-order approximation showing that each farm’s conditional loss—originally expressed through its own Generalized Linear Model (GLM)—can be closely approximated by a single common weather index plus a correction term. This result bridges heterogeneous local loss models and a global reference index, making parametric insurance both fair and tractable. To turn theory into practice, we develop a computational framework that estimates the necessary conditional moments via non-parametric kernel methods and enriches limited weather data through copula-based simulation.
On the empirical side, we apply our hybrid model to 50 solar farms in southern Germany. By combining physics-based day-ahead forecasts with high-resolution climate data, we compute each farm’s revenue shortfall, build the weather index at a reference site, and layer on a P2P basis risk sharing mechanism. In this real-world setting, our approach reduces loss volatility by 55%, demonstrating its effectiveness in stabilizing returns and strengthening the financial case for renewable energy investments.
@@
2 A combined parametric insurance and P2P basis risk sharing mechanism
This section introduces the combined parametric insurance and P2P basis risk sharing mechanism.
@@ 2.1.1 Individual loss and basis risk
In renewable energy production, the random loss $X_{i}$ for producer $i$ is the financial loss associated with forecast uncertainties at a given date. The parametric insurance is triggered when an observed index $Z$ based on weather-related variables for a reference location reaches a predefined threshold $z_{0}$. Like the Consumer Price Index (CPI), which tracks changes in the cost of a standard basket of goods and services over time, the weather index $Z$ quantifies weather risks for renewable energy production through a composite of weather variables. An increase in the index reflects a rise in extreme weather events, leading to very low production, highlighting elevated risk levels for renewable energy systems.
@@ 2.1.2 Risk pooling and sharing
where $\sigma_{i}^{2}(z)=\text{Var}[X_{i}\mid Z=z],$ is the conditional variance of $X_{i}$ for producer $i$ given that the weather index $Z$ is equal to $z$ that exceeds the threshold $z_{0}$. This formula is not only a proportional allocation based on conditional variances but also serves as an approximation of the conditional expectation $\mathbb{E}[\varepsilon_{i}\mid S_{\varepsilon},Z=z]$ when $n$ is large, as shown in Denuit and Robert [13]. It ensures fairness and aligns each producer’s contribution with the variability of their losses. The derivation of this formula as an approximation of $\mathbb{E}[\varepsilon_{i}\mid S_{\varepsilon},Z=z]$ assumes that the basis risks $\varepsilon_{i}$ are independent, conditional on $Z$. This conditional independence is crucial, as it requires the weather index $Z$ to act as a sufficiently informative linear score, capturing all potential conditional dependencies between individual losses. $Z$ should comprehensively reflect the shared weather-related risk factors influencing producers’ financial outcomes. By carefully designing $Z$, this assumption becomes a valid and practical simplification, ensuring the effectiveness and fairness of the risk-sharing mechanism.
@@ 2.2 Parameterizing the insurer’s index benchmark for effective risk sharing
In this paper, the construction of the index $Z$ draws inspiration from the North American Actuaries Climate Index (ACI) in the sense that both use a linear combination of climate variables (Garrido et al. [18]). While the methodology used to build our index differs from the ACI, this linear structure is standard across many actuarial climate indices, providing an intuitive way to summarize weather conditions relevant to risk assessment and insurance design.
Remark 1.
@@
We now detail the approximation of the following quantities of interest: $\mathbb{E}\left[X_{i}\mid Z=z\right]$ and $\text{Var}\left[X_{i}\mid Z=z\right]$. The major theoretical issue in this paper lies in the fact that the conditional expectation $\mathbb{E}[X_{i}\mid\bm{a}_{i}^{\prime}\bm{Y}]$ is known analytically (because it is deduced from an individual GLM), but not $\mathbb{E}[X_{i}\mid Z]$. However, it is possible to find an approximation which depends only on the density distribution function of $Z$ and moments of the distribution $\bm{Y}$ given $Z$. These moments, even if they do not have an analytical expression, can be obtained using a non-parametric estimation method based on simulations.
@@ 2.2.1 Approximation of the conditional expectation $\mathbb{E}\left[X_{i}\mid Z=z\right]$
Proposition 1 (Second-order approximation of conditional expectation).
Let us assume that $\bm{a}$ and $\bm{a}_{i}$ are not collinear vectors and that $\left\|\bm{a}_{i}-\bm{a}\right\|$ is small. There exist two functions $L_{i}$ and $F_{i}$ such that:
@@
The proof of the approximation of the conditional expectation of the loss in Proposition 5 is given in Appendix B. This result formalizes the idea that, when the losses $X_{i}$ depend on a vector of covariates $\bm{Y}$, their conditional expectation given a linear score $\bm{a}_{i}^{\prime}\bm{Y}$ can be approximated by the conditional expectation given a common reference index $Z=\bm{a}^{\prime}\bm{Y}$, up to a second-order correction. This is crucial in our context because:
  • The theoretical conditional expectation $\mathbb{E}[X_{i}\mid Z=z]$ is required for basis risk sharing computations, but it is not directly estimable from the GLM.
  • Each park $i$ has its own weight $\bm{a}_{i}$, leading to model heterogeneity. However, operational implementation of the index requires a common threshold based on a shared index $Z$.
This approximation thus bridges individual heterogeneity and a pooled reference structure, and is key to making the parametric index tractable for real-world insurance design.
To compute $\mathbb{E}[X_{i}\mid Z=z]$ in Proposition 5, we develop a computational approach that estimates the relevant conditional expectations and densities using kernel-based methods. Because the approximation involves quantities that are not analytically tractable, this non-parametric strategy is essential. To address the challenges of limited data, we simulate additional weather data by calibrating the marginal distributions of the weather variables and modeling their dependencies with a vine copula [19]. This enriched dataset improves the stability and reliability of the conditional expectation estimation. Further implementation details are provided in Appendix C and support the empirical analysis in Section 3.
@@ 2.2.3 Final optimisation
3 Case study: physics-based simulation and P2P basis risk sharing in southern Germany
@@ 3.1.1 Power plants data
We compiled a dataset of large-scale solar power plants in Germany using the German governmental database ; see, e.g., Bundesnetzagentur [20]. This dataset was constructed by extracting and processing information on operational solar farms, focusing on installations exceeding 5 megawatt (MW). Such large-scale projects represent potential candidates for parametric insurance, as they primarily operate within electricity markets rather than for self-consumption.
@@ 3.1.2 Recovery of the energy production loss
To quantify these imbalances, we reconstruct both forecasted and realized conditions using meteorological variables from two datasets provided by the European Centre for Medium-Range Weather Forecasts (ECMWF). The historical weather variables are obtained from the ERA5 reanalysis dataset [21], which provides global coverage at a spatial resolution of 31 km with an hourly temporal resolution. This dataset, widely used in renewable energy research, spans 2012 to 2022 and includes near-surface air temperature (T2m) and surface solar radiation downwards (SSRD), which are essential for forecasting solar production.
@@
For each given location, we model both forecasted and realized solar electricity production using the PVlib library [22] in Python, which simulates photovoltaic system performance based on climate variables. The energy production loss for each solar farm is then defined as the difference between its forecasted and actual production. The PVlib model relies on several key inputs, including SSRD, T2m, panel orientation, and panel tilt angle. The tilt angle, which represents the inclination of the solar panels relative to the horizontal plane, is set at 25° in our simulations. This value corresponds to the mean tilt across our dataset, where individual panel angles range from 20° to 30°.
To recover energy output, we first compute Direct Normal Irradiance (DNI) and Diffuse Horizontal Irradiance (DHI) from SSRD using the Erbs model [23]. Since these values are not directly available, they must be derived using standard transposition models. The SSRD is then converted into Plane of Array (POA) irradiance, representing the amount of solar radiation incident on the tilted panel surface. The module temperature is estimated using a thermal model that accounts for both air temperature and POA irradiance. Finally, the PVWatts model [24] is used to compute direct current (DC) power, which is then converted into alternating current (AC) power using an inverter model with default efficiency parameters.